OpenStudy (anonymous):

Which of the following is in simplest form? Choose one answer. a. fifteen-twentieths b. four thirty-seconds c. eighteen sixty-sixths d. twenty-seven fortieths

OpenStudy (mathstudent55):

Math is done with numbers, not the spelling of the names of the numbers. \( \dfrac{15}{20}\) \(\dfrac{4}{32} \) \(\dfrac{18}{66} \) \(\dfrac{27}{40} \)

OpenStudy (mathstudent55):

Look at each fraction. Can you divide the numerator and denominator by a number (other than 1)? For example, in the first fraction, 15 and 20 are both divisible by 5, so it's not in simplest form. Now look at each one of the other 3 fractions and see if the numerator and denominator are divisible by the same number.

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

D

OpenStudy (mathstudent55):

4 is divisible by 4 32 is also divisible by 4 That is not it. \( \dfrac{4}{32} = \dfrac{1}{8} \)